OpenStudy (anonymous):

write the equation in logarithmic form 4^7=16234 will award metal!!!!

OpenStudy (jdoe0001):

\(\bf \large a^y=x\implies log_ax\quad thus \quad 4^7=16234\implies log_4\square =\square\)

OpenStudy (anonymous):

\[\log_{4}(16234)=7 \]

OpenStudy (anonymous):

is that the answer?

OpenStudy (jdoe0001):

\(\bf \large a^y=x\implies log_ax=y\quad thus \quad 4^7=16234\implies log_4\square =\square \)